Prove that between any two real roots of the equation
2e^2xsin3x-3=0

There is at least one real root of the equation
e^2xcos3x+1=0
1
Expert's answer
2020-03-02T08:55:05-0500

We consider the following function:

"f(x)=\frac{1}{3}\sin 3x -\frac{1}2{}e^{-2x}"


"\sin 3x" and "e^{-2x}" are continuous and differentiable for all real "x" "\Rightarrow \ f(x)" is continuous and differentiable for all real "x."


Suppose, that "a" and "b" ("a<b)" are real roots of "2e^{2x}\sin 3x-3=0" .

"2e^{2x}\sin 3x-3=6e^{2x}(\frac{1}{3}\sin3x-\frac{1}{2} e^{-2x})=0"

"e^{2x}""\ 0 \Rightarrow \frac{1}{3}\sin3x-\frac{1}{2}e^{-2x}=0."


So, if "a" and "b" are roots of "2e^{2x}\sin 3x-3=0" , then "a" and "b" are roots of "f(x)=0" ,

i.e. "f(a)=f(b)=0."


Now we can use Rolle’s Theorem:

"f(x)" is continuous in "[a,b]" , differentiable in "(a,b)" , "\ f(a)=f(b)=0."

Therefore, "\exists \ c\in(a,b): \ f^\prime (c)=0" .

"f^\prime (c)=\cos(3c)+e^{-2c}=0 \ \Rightarrow e^{2c}\cos3c+1=0"


So, for all real roots ("a" and "b)" of "2e^{2x}\sin 3x-3=0" , there is "c\in(a,b) : e^{2c}\cos3c+1=0" .